According to one survey in India, 75% of Instagram users love REELS. Suppose that 25 Instagram users (randomly selected) have been approached in the university located in vile parle. They have been asked about their status of like/ dislike the Instagram- REELS. a) What is the probability that Exactly 15 of them would agree with the claim (or said they love Insta-REELS)?

b) What is the probability that Exactly 20 of them would agree with the claim (or said they love Insta-REELS)?


1
Expert's answer
2022-03-30T07:28:03-0400

0.75*20= 15

The probability is given by 15/20 = 3/4.